OpenStudy (anonymous):

how do i solve this? for f(x) = 5x+ 25, and g(x) = x +5, find the quotient of f and g.

OpenStudy (anonymous):

factorisation will come in handy here.

OpenStudy (anonymous):

f(x)/g(x) = (5x+25)/(x+5) = (x+5)*5/(x+5) = 5

OpenStudy (anonymous):

The x+5 cancels out in the numerator and denominator

OpenStudy (anonymous):

so i factor the 5x + 25 into (x+5) times 5, and then cancel the x+5 in the numerator and the denominator?

OpenStudy (anonymous):

Yes, correct

OpenStudy (anonymous):

thanks